# How do language models learn facts? Dynamics, curricula and hallucinations

### v1

@@ 4.2 A large portion of the pre-training knowledge is erased during early fine-tuning
Given this uncalibrated starting point, fine-tuning on new individuals is expected to be challenging. We observe a rapid collapse in performance on pre-training data during the initial stages of fine-tuning (first few hundred steps), with very little corresponding acquisition of new knowledge (Figure 5, middle). A larger number of fine-tuning individuals intensifies this effect. Extended fine-tuning eventually changes this trend, and the model starts learning about new individuals. Incorporating replay data about pre-training individuals in the fine-tuning set partially mitigates the initial collapse and facilitates the restoration of corrupted knowledge (Figure 5, right). Finally, we do not observe such degradation when fine-tuning on subsets of the pre-training data (see the corresponding analysis in Appendix F.3). We complement our analysis of fine-tuning dynamics with one in which the training distribution changes regularly (Appendix F.5). These experiments confirm the findings reported here, as well as highlight that it becomes easier to learn new individuals as training progresses.
We now turn to explaining this behavior. The first hypothesis we consider is related to attention patterns. Indeed, introducing new individuals may disrupt attention patterns, requiring time to restore recall ability, while keeping parametric knowledge relatively untouched, and further training may restore them. We analyze the network’s attention patterns, in both our fine-tuning (Figure O) and switching distribution (Figure X) setups, and find them to be remarkably stable, therefore mostly ruling out this hypothesis. Our second hypothesis is that introducing novel individuals creates additional key-value pairs that corrupt existing ones when being learned. To test this hypothesis, we train a one hidden layer multi-layer perceptron on a toy associative recall task (Appendix F.4). With such a model, which has no attention layers at all, we are able to reproduce the main qualitative findings of the experiments above, suggesting that changes in the feed-forward associative memories of the model we consider are the main factor explaining the model’s behavior. Getting a finer understanding of these dynamics constitute an interesting future research direction. For example, investigating how introducing independent knowledge topics affect this early catastrophic forgetting could provide some practically relevant insights.
@@ B.1 Rationale behind our design choices
We build upon the synthetic biography dataset and analysis of Allen-Zhu and Li (2023), as well as the mechanistic interpretability studies of Geva et al. (2023) and Nanda et al. (2023b). Geva et al. (2023) and Nanda et al. (2023b) find that pretrained large language models encode individual-specific information in their residual stream as soon as the name of the individual is encountered. Allen-Zhu and Li (2023) observe such a behavior when training models on biographies whose order is permuted every new occurrence, along with having multiple biographies for “celebrities” (see their Q-probing analysis). We hypothesize that textual variation within the biography distribution is key to achieve this behavior. Therefore, our setup ensures unique biographies by introducing 25 distinct templates per attribute type (see next section) and permuting their order. Importantly, biographies are resampled for each new sequence. Figure A demonstrates the critical importance of random ordering and the need for some textual diversity within individual biographies. While we have not ablated the total number of templates per attribute type, we believe that the overall textual diversity of the distribution is necessary for the learning of robust features. We always repeat the individual’s name in every sentence. In our experiments, the individual’s name is repeated in every sentence to facilitate knowledge development, though we believe later occurrences could be replaced with pronouns without significantly affecting results, as (Allen-Zhu and Li, 2023) found the benefits of full name repetition primarily in the non-permuted and not so diverse regime.
@@ B.2 Details about the biography generation process
  • Full name. First and middle names are sampled from a list of 900, and last names from a list of 1,000, resulting in 810 million potential unique names. We ensure that each individual has a unique full name. These names are among the most used worldwide.

### v2

@@ 4.2 A large portion of the pre-training knowledge is erased during early fine-tuning
Given this uncalibrated starting point, fine-tuning on new individuals is expected to be challenging. We observe a rapid collapse in performance on pre-training data during the initial stages of fine-tuning (first few hundred steps), with very little corresponding acquisition of new knowledge (Figure 5, middle). A larger number of fine-tuning individuals intensifies this effect. Extended fine-tuning eventually changes this trend, and the model starts learning about new individuals. Incorporating replay data about pre-training individuals in the fine-tuning set partially mitigates the initial collapse and facilitates the restoration of corrupted knowledge (Figure 5, right). Finally, we do not observe such degradation when fine-tuning on subsets of the pre-training data (see the corresponding analysis in Appendix F.3). These results are consistent with the findings of Gekhman et al. (2024), who found that fine-tuning on new knowledge is slow and leads to hallucinations. We complement our analysis of fine-tuning dynamics with one in which the training distribution changes regularly (Appendix F.5). These experiments confirm the findings reported here, as well as highlight that it becomes easier to learn new individuals as training progresses.
We now turn to explaining this behavior. The first hypothesis we consider is related to attention patterns. Indeed, introducing new individuals may disrupt attention patterns, requiring time to restore recall ability, while keeping parametric knowledge relatively untouched, and further training may restore them. We analyze the network’s attention patterns, in both our fine-tuning (Figure O) and switching distribution (Figure X) setups, and find them to be remarkably stable, therefore mostly ruling out this hypothesis. Our second hypothesis is that introducing novel individuals creates additional key-value pairs that corrupt existing ones when being learned. To test this hypothesis, we train a one hidden layer multi-layer perceptron on a toy associative recall task (Appendix F.4). With such a model, which has no attention layers at all, we are able to reproduce the main qualitative findings of the experiments above, suggesting that changes in the feed-forward associative memories of the model we consider are the main factor explaining the model’s behavior. Getting a finer understanding of these dynamics constitutes an interesting future research direction. For example, investigating how introducing independent knowledge topics affects this early catastrophic forgetting could provide some practically relevant insights.
@@ 5 Discussion
Limitations
In our study, we train relatively small language models on a synthetic factual recall task to investigate knowledge acquisition dynamics. While this approach enables precise experimental control and mechanistic insights, it presents several important limitations.
Scale and architecture.
The $44$M parameter models we train are substantially smaller than the billion-parameter models used in practice. The learning dynamics we observe may manifest differently at larger scales or with different architectural choices. However, our ablations across model sizes, architectures (including recurrent variants), and hyperparameters suggest the core phenomena are robust to these variations.
Data.
The synthetic factual recall task, while designed to capture essential properties of knowledge acquisition, only represents one specific type of knowledge language models acquire during pretraining. Different types of knowledge could interact in sophisticated ways that are not modeled in our task. This may create extra redundancy that language models could leverage to accelerate learning. Nevertheless, our fine-tuning results already align with practically observed inefficiencies of fine-tuning for integrating new knowledge, suggesting some degree of transferability.
Despite these limitations, we believe our work provides a necessary foundation for understanding knowledge acquisition in neural language models. The mechanistic insights we derive offer concrete hypotheses that future work can validate in more realistic scenarios. We hope these findings will guide the design of more effective training strategies and data scheduling approaches for large-scale language models.
@@ B.1 Rationale behind our design choices
We build upon the synthetic biography dataset and analysis of Allen-Zhu and Li (2023), as well as the mechanistic interpretability studies of Geva et al. (2023) and Nanda et al. (2023b). Geva et al. (2023) and Nanda et al. (2023b) find that pretrained large language models encode individual-specific information in their residual stream as soon as the name of the individual is encountered. Allen-Zhu and Li (2023) observe such a behavior when training models on biographies whose order is permuted every new occurrence, along with having multiple biographies for “celebrities” (see their Q-probing analysis). We hypothesize that textual variation within the biography distribution is key to achieve this behavior. Therefore, our setup ensures unique biographies by introducing 25 distinct templates per attribute type (see next section) and permuting their order. Importantly, biographies are resampled for each new sequence. Figure A demonstrates the critical importance of random ordering and the need for some textual diversity within individual biographies. While we have not ablated the total number of templates per attribute type, we believe that the overall textual diversity of the distribution is necessary for the learning of robust features. In our experiments, the individual’s name is repeated in every sentence to facilitate knowledge development, though we believe later occurrences could be replaced with pronouns without significantly affecting results, as (Allen-Zhu and Li, 2023) found the benefits of full name repetition primarily in the non-permuted and not so diverse regime.
@@ B.2 Details about the biography generation process
  • Full name. First and middle names are sampled from a list of 900, and last names from a list of 1,000, resulting in 810 million potential unique names. We ensure that each individual has a unique full name. These names are among the most commonly used worldwide.